### v2

# A Universal Relation Between Intermittency and Dissipation in Turbulence

Abstract
The relationship between fundamental quantities of turbulent flows like the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$, expected to be constant for homogeneous isotropic turbulence, is explored with respect to their relationship for a wider class of non-ideal turbulent flows. For the energy cascade, $C_{\varepsilon}$ is known to reflect its basic overall properties, while $\mu$ quantifies the emerging intermittency down the cascade. Using a very large dataset of turbulent wakes, grid-generated turbulence and an axisymmetric jet, we find that $\mu$ is inversely proportional to $C_{\varepsilon}$, providing a new empirical principle.
Turbulence is still one of the major unsolved problems in physics [1]. The lack of a theory of turbulent flows hinders our modelling capabilities in several fields, as there are many situations where turbulence has a crucial influence, with turbulent wakes generated by wind turbines being currently among the most relevant industry-related applications [2, 3]. The most common approach to tackle this problem is to define equations where factors or exponents are obtained by dimensional arguments or experiments. Within this approach, a relevant question that arises concerns the universality of these constants [4, 5]. Since the publication of Kolmogorov’s phenomenology in 1941 (K41) [6], there has been intense work to shed light on this question. This was enhanced by the publication of Kolmogorov’s revised theory in 1962 [7], as further constants appeared. Although there were some deviations in experimental results, the overall opinion – at least until the 1990s – was that such constants, for sufficiently large Reynolds numbers, are universal for fully developed turbulence, which led to a focus on their numerical values [4, 8, 9, 10, 11].
Among all relevant parameters, two of the most important constants to describe turbulent flows are, arguably, the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$. The former has been identified as one of the most important parameters to describe turbulence [12], as it can be directly linked to the energy cascade [13]. On the other hand, the constant $\mu$, first introduced by Kolmogorov’s 1962 theory [7], and later redefined by Castaing and collaborators [14, 9], quantifies departures from self-similarity within the inertial range. Despite its relevance, a physical interpretation of $\mu$ is still missing. Furthermore, over the last two decades, the universality of $C_{\varepsilon}$ has been challenged by several experimental and numerical studies [13].
The dissipation constant is defined as
$C_{\varepsilon}=\frac{\varepsilon\>L}{u^{\prime 3}},$ (1)
with $\varepsilon$ the mean turbulent kinetic energy dissipation rate, $u^{\prime}$ the RMS value of the streamwise fluctuating velocity and $L$ the integral length scale. The importance of $C_{\varepsilon}$ for modelling arises from the fact that it gives an estimation of the mean dissipation rate, which is often needed to have a closed system of equations [15]. According to K41 phenomenology, considering stationary, homogeneous turbulence, $C_{\varepsilon}$ should depend on boundary conditions, being constant for sufficiently large values of the Taylor-length based Reynolds number $Re_{\lambda}=u^{\prime}\,\lambda/\nu$, with $\lambda$ the Taylor length scale and $\nu$ the kinematic viscosity of the fluid.
In several free-shear flows, results not consistent with K41 phenomenology have been reported, and the value of $C_{\varepsilon}$ at the centreline has been found to decrease with increasing $Re_{\lambda}$ [13, 16]. More precisely, for this so-called non-equilibrium cascade, $C_{\varepsilon}$ is expected to be linear with $\sqrt{Re_{G}}/Re_{\lambda}$ (with $Re_{G}$ a Reynolds number that depends on the inlet properties of the flow, such as the wind tunnel inflow velocity and the diameter of the wake-generating object). Most of the work on this topic aims at quantifying the dependence of $C_{\varepsilon}$ over a broad range of Reynolds numbers. These studies concern laminar inflows, and little attention has been paid to further details of the energy cascade, such as emerging intermittency.
The intermittency factor $\mu$ is used to characterise deviations from K41 and is a measure of how “fast” the cascade propagates to smaller scales. This deviation is seen in the changing statistics of the averaged velocity increments $\delta_{u}(r):=u(x+r)-u(x)$ ($r$ being the spatial increments), which can be quantified using the scale-dependent shape factor $\Lambda^{2}(r)$. $\Lambda^{2}$, that is linked to the dispersion of the logarithm of the – also scale-dependent – energy dissipation [7], can be estimated as [17],
$\Lambda^{2}(r)=\frac{\ln\left(F(\delta_{u}(r))/3\right)}{4},$ (2)
where $F(\delta_{u}(r))$ denotes the flatness. In the inertial range, for $\eta\ll r\ll L$ (with $\eta=(\nu^{3}/\varepsilon)^{\frac{1}{4}}$, being the Kolmogorov length scale), the scale dependence is expected to follow the relation [7, 14],
${\Lambda^{2}=\Lambda_{0}^{2}+\mu/9\;\ln(L/r),}$ (3)
with $\Lambda_{0}^{2}$ a constant that depends on the properties of the flow (see also figure 1). Note, that the relation (3) is consistent with the power-law scaling behavior of the energy spectral density in the inertial range, as discussed below (see equation 4).
Within this rationale, the non-Gaussianity of the velocity increment distributions can be quantified by a single scalar parameter. The value of $\mu$ has been commonly accepted to be constant at large values of the $Re_{\lambda}$. The reported values range between $0.2$ and $0.4$ for homogeneous isotropic turbulence (HIT) and a value of $0.26$ has been established as representative of turbulent flows [9, 18, 3].
Despite intense research on turbulent flows, no relation between $\mu$ and $C_{\varepsilon}$ has been predicted nor found. Furthermore, the universality of these constants remains an open question, with a growing amount of evidence pointing against this feature. While most studies deal with homogeneous turbulence, in this study we also focus on inhomogeneous flows. Hence, we present a detailed experimental work on turbulent wakes supported by previously published datasets on grid-generated turbulence and a turbulent axisymmetric jet.
Figure 1: Schematic representation of the spatial expansion of the turbulence cascade: the blue curve shows the energy spectrum, with energy transported to smaller scales, while the red curve illustrates the evolution of the shape factor $\Lambda^{2}(r)$ for normalised velocity increments.
We carried out wind tunnel experiments in LEGI, Grenoble by generating a turbulent wake behind porous and impermeable objects under different inflow conditions. The facility is a closed-circuit system with test section dimensions of 0.75 $\times$ 0.75 $\times$ 444\,4 m3 (see supplementary material [19]). The objects were cylinders and disks while both the diameter and solidity vary (see also table I in the supplementary material [19]). To reduce the blockage, the cylinders were directly screwed to the walls of the tunnel while the disks were hold at the tunnel’s centreline with four thin piano wires which went through taped holes in the tunnel walls. For varying the inflow, both the entrance of the test section and the tunnel freestream velocity were modified. We used laminar inflow, and different turbulent background flows generated by a static regular grid with cylindrical bars or an active grid driven in a triple random mode (information about the turbulent flow produced by grids is reported in a previous work [20]). The wake of the objects was scanned by a 111\,1D-hot-wire with a temporal resolution of 505050\,50kHz (with an anti-aliasing filter set at 303030\,30kHz) which was traversed in streamwise and spanwise directions. The conversion from the temporal to the spatial domain was made using Taylor’s hypothesis. The nearest streamwise position to the wake generator was $5\,D$ and the farthest was $200\,D$, with $D$ being the diameter of the object. For every position the data was acquired for 120120120\,120s. In total, $15$ configurations were carried out and $505$ centreline points were taken. For simplicity, only points in centreline are considered 111pre-analysis shows that our results also hold for non-centreline data. Additionally, the background flows without the object were also measured (see table I in the supplementary material [19] for further information.)
As discussed above, we complement our study with $113$ velocity time series from three previously published experimental datasets covering different canonical flows [22, 23, 24] (for further details see supplementary material [19]). Within those measurement configurations, five involve decaying grid-generated turbulence and one captures a free axisymmetric jet. All cases use air as the fluid and the velocity time signals were collected by means of a 111\,1D-hot-wire.
For all those centreline points from all mentioned datasets, a rigorous analysis of common 111\,1D-free-shear turbulence constants is derived and checked individually. Only time series that have the following characteristics were kept for further analysis:
a) the turbulence intensity $TI=u^{\prime}/\overline{u}$ with $\overline{u}$ being the actual mean velocity, decreases in streamwise direction, b) the large scale increments have an almost Gaussian distribution ($\Lambda_{0}^{2}<0.025$) c) the slope of the energy spectrum within the inertial range is smaller than $-1.25$, d) the length of the inertial range for the energy spectrum is larger than $0.5$ decades, e) the $R^{2}$-value of the linear regression of the fit of $\Lambda^{2}$ is larger than $0.99$, f) the mean squared deviation from the calculation of $\Lambda^{2}$ is smaller than $0.11$, g) the flatness of the velocity time signal is around $3$ ($3.12.67<F<$ 3.1) and h) the skewness of the time signal is below $0.1$. Details and an example of the analysis is shown and discussed in the supplementary material [19]. After this conditioning, 313 velocity time series remained for the analysis. Each of them represents at least $2\times 10^{4}$ integral time scales.
The integral length scale $L$ is calculated by integrating the autocorrelation function until it crosses $1/\mathrm{e}$ ($\mathrm{e}$ being the Eulerian number). Other integration limits were tested, and the trends presented in this Letter remained unchanged. $\varepsilon$ is derived – the common definition for HIT conditions is used – by integrating the dissipation spectrum $\varepsilon=\int_{0}^{\infty}15\nu\>k^{2}E(k)\>\mathrm{d}k$ (with $k$ the wavenumber in the Fourier space) and modelling the high-frequency noise with an exponential fit (a test with different fitting limits and functions gave an estimated error in the estimation of $\varepsilon$ in the order of 5%). We extracted, after equation (3), $\mu$ by fitting the evolution of the shape factor $\Lambda^{2}$ within the inertial range [14].
First, we investigate if the parameters $\mu$ and $C_{\varepsilon}$ present a constant value in our dataset, including all the flows studied. In figure 2a) it is clear that $\mu$ presents a large spreading and it does not approach towards an asymptotic limit at large values of $Re_{\lambda}$. (For reference, the dependency of $\mu$ with the streamwise distance to the turbulence generator is shown in the supplementary material [19].) Figure 2b) shows that $C_{\varepsilon}$ also behaves similarly, i.e. $C_{\varepsilon}$ does not have a constant value.
Having a closer look on the data we see that for each set of boundary conditions, a linear relation between $C_{\varepsilon}$ and $\sqrt{Re_{G}}/Re_{\lambda}$ is consistent with the dissipation scalings from non-equilibrium turbulence [13]. Remarkably, when such conditions are changed, the linear relation still applies but with a different slope, showing the emergence of new length scales that affect the dissipation scalings.
Overall, our dataset does not show a universal or predictable behaviour in terms of neither $C_{\varepsilon}$ nor $\mu$. Next, we want to see if there is a relation between $\mu$ and $C_{\varepsilon}$. Figure 3a) shows that $\mu\propto 1/C_{\varepsilon}$, or, respectively, that the product of those two quantities is constant. In order to quantify this, initially a simple fit is presented through the data points. Note that the commonly accepted value of $\mu$ of around 0.26 is not the asymptotic limit of our measurements. Moreover, the product of $\mu$ and $C_{\varepsilon}$, here defined as $\alpha$, is shown as a function of $Re_{\lambda}$ in figure 3b). Remarkably, $\alpha$ not only remains approximately constant for all cases, it also does not show any trends with $Re_{\lambda}$. In the supplementary material we provide further figures and a table showing that the RMS value of $\alpha$ remains lower than the individual corresponding RMS values of $C_{\varepsilon}$ and $\mu$ [19]. To the authors best knowledge, this is the first time a clear relation between these constants is found.
[…]
Figure 3: a) $\mu$ as a function of $C_{\varepsilon}$ for the highly restricted data. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence [9], and the black line corresponds to a least-square fit ($R^{2}=0.86$ with the fitted values of $0.1067\pm 0.0025$ and $0.0227\pm 0.0091$). b) $\alpha$ versus $Re_{\lambda}$ for the highly restricted data. The red line indicates the result for $\alpha$ from the fit from a) while the black solid line represents the actual mean value of the ensemble of $\alpha$ values. Additionally, two black dashed lines indicate the corresponding standard deviations from the mean. For both figures, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the supplementary material [19].
We continue by investigating how two other relevant parameters behave, namely the Kolmogorov constant $C_{k}$ and the power law governing the energy spectral density within the inertial range. Following the notation of  [25],
${E(k)=C_{k}\>\>\varepsilon^{2/3}\>k^{-5/3}\>(k\>\eta)^{-\gamma+5/3},}$ (4)
[…]
In figure 4a) & b), the dependency of $\gamma$ on $C_{\varepsilon}$ as well as $C_{k}$ on $\gamma$ are presented, respectively ($C_{k}$ as a function of $C_{\varepsilon}$ is shown in the supplementary material [19]). Interestingly, the relation between $\mu$ and $C_{\varepsilon}$ is transferred to the other quantities. Moreover, all quantities present large variations with respect to the expected values for HIT within the K41 phenomenology. We remark that our results are still consistent with those reported in other experiments in inhomogeneous flows, such as other turbulent wakes [3] or atmospheric turbulence [27].
Figure 4: a) $\gamma$ as a function of $C_{\varepsilon}$ and b) $C_{k}$ versus $\gamma$ for the highly restricted data. The red lines indicate both the commonly accepted value for $C_{k}$ and $\gamma$ for HIT [8], while the black lines correspond to a least-square fit for a) and b) with $R^{2}$ being $0.8$ and $0.98$, respectively. Once more, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the supplementary material [19].
To conclude, based on the study of a very large number of velocity time series, the joint and individual behaviour of the dissipation constant and the intermittency factor were quantified, as well as the spectral law of turbulence. In contrast to the common approach of more restrictive data selection to obtain more precise results on turbulence, we adopted a less restrictive data selection method. This follows the idea of Vassilicos [13], which allows for the investigation of states of non-equilibrium turbulence. For the different canonical turbulent flows that compose our dataset, our analysis leads to the proposition of a new relation,
$\mu\>C_{\varepsilon}=\alpha=\text{const.},$. , (5)
replacing the previous proposition that these quantities should be individually constant.
Figure 3b indicates that there is no Reynolds number dependence of the product $\mu\>C_{\varepsilon}$. This is a remarkable result. It has long been claimed that certain turbulent properties are universal, i.e., independent of the type of flow, for large enough Reynolds numbers. Together with the relations to other turbulent constants (shown in figure 4), this new proposed relation combines the dimensionless energy dissipation as an overall quantity of turbulent states with the statistics of velocity fluctuations at two points. Note that the power spectrum and the intermittency factor fully capture the two-point statistics [29]. Thus, we observe a new universality combining overall constants ($C_{\varepsilon}$ or $C_{k}$) with the two-point statistics of the velocity field.
As a remark, it should be pointed out that our results do not imply the absence of Reynolds number effects. For example, the depth of the cascade is still expected to depend on $Re_{\lambda}$, influencing the range of the power law of the power spectrum and therefore the intermittency at the smallest scales (which are also expected to be determined by the Reynolds number). We also observe that the evolution of $C_{\varepsilon}$ with the Reynolds number follows different linear relations sections for different flows, as discussed above.
Next, we set our result in the context of the $\Pi$-theorem of Vaschy-Buckingham [30, 31]. $C_{\varepsilon}$ and $\mu$ can also be viewed as two $\Pi$-parameters for dissipation and the scale dependency of the fluctuations of dissipation. For different turbulent flow configurations, it is a challenge to find the functional relation between these $\Pi$-parameters. For a prominent example, see the Rayleigh-Bénard convection [32]. In this sense, our result reads as $\mu\>C_{\varepsilon}=\Pi_{1}\cdot\Pi_{2}=\text{const.}\neq f(Re)$.
Further, a relevant consequence of equation (5) is that we can define $\mu$ as,
$\mu=\frac{u^{\prime 3}\>\alpha}{\varepsilon\>L}.$ (6)
This relation has the advantage that the quantities on the right-hand side are much easier to measure and converge than $\mu$ (when estimated using standard approaches).
While our results appear to be applicable to a large variety of flows, further work is needed to assess whether our findings apply to other inhomogeneous and unsteady flows, which may lead to an even greater validity of our new law for turbulent flows.
Acknowledgements.
This work has been supported by a LabEx Tec21 grant (Investissements d’Avenir – grant agreement no. ANR-11-LABX-0030). We thank Christophe Penisson, Muriel Lagauzere, Stephane Pioz-Marchand and Sylvain Dauge for helping with the experiment, Valentin Groß for proof-reading, Olivier De Marchi and Sebastian Bergemann for the IT support, Thomas Messmer, Lars Neuhaus and Daniela Moreno for validating parts of the used code and Michael Hölling, Jan Friedrich and Christos Vassilicos for fruitful discussion.