### v3

# A universal relation between intermittency and dissipation in turbulence

Abstract
Fundamental quantities of turbulent flows, such as the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$, are examined in relation to each other for a broader class of non-ideal turbulent flows. In the context of the energy cascade, it is known that $C_{\varepsilon}$ reflects its basic overall properties, while $\mu$ quantifies the intermittency that emerges throughout the cascade. Using an extensive hot-wire dataset of turbulent wakes, grid-generated turbulence, and an axisymmetric jet, we individually analyze these quantities as one-dimensional surrogates of the energy cascade, considering only data that exhibit consistent scaling behavior. We find that $\mu$ is inversely proportional to $C_{\varepsilon}$, offering a new empirical principle that bridges the gap between large and small scales in arbitrary turbulent flows.
preprint: AIP/123-QED
Despite recent advances in the study of turbulent flows sreenivasan2025turbulence, solving the Navier–Stokes equations at large Reynolds numbers and for an arbitrary set of boundary conditions still remains beyond current numerical and theoretical capabilities frisch1995turbulence; alberti2023still. The lack of a simplifying framework for complex turbulent flows hinders modeling efforts across various fields, as there are many engineering problems where turbulence plays a crucial role, turbulent wakes generated by wind turbines being among the most relevant industry-related applications stevens2017flow; neunaber2020distinct. The most common approach to tackle turbulence is to define equations in which factors or exponents are determined by dimensional arguments or experiments. Within this approach, a relevant question that arises concerns the universality of these constants, even when considering homogenous isotropic turbulence (HIT) only praskovsky1994measurements; renner2002universality.
Since the publication of Kolmogorov’s phenomenology in 1941 (K41) kolmogorov1941local, there has been intense work to shed light on this question. This was enhanced by the publication of Kolmogorov’s revised theory in 1962 kolmogorov1962refinement and further developments like the multifractal formalism (cf. sreenivasan1991fractals), as further constants appeared. Although there were some deviations in experimental results, the overall opinion – at least until the 1990s – was that such constants, for sufficiently large Reynolds numbers, are universal for fully developed turbulence, which led to a focus on their numerical values praskovsky1994measurements; sreenivasan1995universality; arneodo1996structure; praskovsky1997comprehensive; yeung1997universality.
Among all relevant parameters, two of the most important constants to describe turbulent flows are, arguably, the dissipation constant $C_{\varepsilon}$ and the intermittency factor $\mu$. The former has been identified as one of the most important parameters to describe turbulence lumley1992some, as it can be directly linked to the energy cascade vassilicos2015dissipation. On the other hand, the constant $\mu$, first introduced by Kolmogorov’s 1962 theory kolmogorov1962refinement, and later redefined by Castaing and collaborators castaing1990velocity; arneodo1996structure, quantifies departures from self-similarity within the inertial range. Despite its relevance, a physical interpretation of $\mu$ is still missing. Nevertheless, and independently on the formalism used to interpret this quantity, it provides a scalar value that quantifies the relevance of intermittency as departures from Gaussianity of velocity derivatives and their increments in the inertial range. Furthermore, over the last two decades, the universality of $C_{\varepsilon}$ (sometimes referred to as “zeroth law” of turbulence sreenivasan2025turbulence) has been challenged by several experimental and numerical studies vassilicos2015dissipation.
The dissipation constant is defined as
$C_{\varepsilon}=\frac{\varepsilon\>L}{u^{\prime 3}},$ (1)
with $\varepsilon$ the mean turbulent kinetic energy dissipation rate, $u^{\prime}$ the RMS value of the streamwise fluctuating velocity and $L$ the integral length scale. The importance of $C_{\varepsilon}$ for modelling arises from the fact that it gives an estimation of the mean dissipation rate, which is often needed to have a closed system of equations george1989self. According to K41 phenomenology, considering stationary, homogeneous turbulence, $C_{\varepsilon}$ should depend on boundary conditions, being constant for sufficiently large values of the Taylor-length based Reynolds number $Re_{\lambda}=u^{\prime}\,\lambda/\nu$, with $\lambda$ the Taylor length scale and $\nu$ the kinematic viscosity of the fluid.
In several free-shear flows, results not consistent with K41 phenomenology have been reported, and the value of $C_{\varepsilon}$ at the centreline has been found to decrease with increasing $Re_{\lambda}$ vassilicos2015dissipation; ortiz2021high. More precisely, for this so-called non-equilibrium cascade, $C_{\varepsilon}$ is expected to be linear with $\sqrt{Re_{G}}/Re_{\lambda}$ (with $Re_{G}$ a Reynolds number that depends on the inlet properties of the flow, such as the wind tunnel inflow velocity and the diameter of the wake-generating object). Most of the work on this topic aims at quantifying the dependence of $C_{\varepsilon}$ over a broad range of Reynolds numbers. These studies concern laminar inflows, and little attention has been paid to further details of the energy cascade, such as emerging intermittency.
Intermittency is one of the causes of deviations from the K41 phenomenology. One of its most significant features is that velocity increments, $u_{r}(x):=u(x+r)-u(x)$, exhibit extreme events at small scales, $r$. There are two standard approaches to quantify intermittency.
The most common one is to investigate the scaling behavior of generalized structure functions of a given order $n$, i.e., the $n$-th order moment of velocity increments, $\overline{(u_{r}(x))^{n}}$, where the over-bar implies a spatial averaging. As K41 implies that $\overline{(u_{r}(x))^{n}}\propto(\varepsilon\>r)^{\zeta_{n}}$, with the exponent $\zeta_{n}$ equal to $n/3$, many works focus on corrections either to the power-law form or to the value of the exponent sinhuber2017dissipative; reinke2018universal, cf. sreenivasan1997phenomenology. This approach has been extensively used for HIT, resulting in various scaling models for the exponent $\zeta_{n}$ of the structure functions. Starting from Kolmogorov’s 1962 theory (based on a lognormal model for the two-point energy dissipation), many improvements and alternative frameworks have been proposed, such as the works of She-Leveque she1994universal, Castaing-Yakhot castaing1996temperature; yakhot1998probability, and L’vov-Procaccia l2000analytic. For a recent review, the reader is referred to Benzi & Toschi benzi2023lectures.
In this structure function approach, two questions arise. On the one hand, as discussed above, an open question is how the higher-order exponents $\zeta_{n}$ deviate from the relation $\zeta_{n}=n/3$ implied by K41. Typically, these deviations become significant for $n>4$, while differences between anomalous scaling models become significant for $n>6$. On the other hand, another open discussion concerns how deviations from the expected scaling of the structure function can be understood in the context of non-isotropic turbulent flows Arad1998; biferale2002probing; kurien2002measures; benzi2023lectures; zhou2021turbulence.
A second approach is to investigate directly the form of the probability density functions $p(u_{r}(x))$ at different values of $r$. This can be done via the scale-dependent shape factor $\Lambda^{2}(r)$. This quantity, that is linked to the dispersion of the logarithm of the – also scale-dependent – energy dissipation kolmogorov1962refinement; castaing1990velocity, can be estimated as chilla1996multiplicative,
$\Lambda^{2}(r)=\frac{\ln\left(F(u_{r}(x))/3\right)}{4},$ (2)
where $F(u_{r}(x))$ denotes the flatness. In the inertial range, for $\eta\ll r\ll L$ (with $\eta=(\nu^{3}/\varepsilon)^{\frac{1}{4}}$, being the Kolmogorov length scale), the scale dependence is expected to follow the relation kolmogorov1962refinement; castaing1990velocity,
$\Lambda^{2}(r)=\Lambda_{0}^{2}+\mu/9\;\ln(L/r),$} (3)
with $\Lambda_{0}^{2}$ a constant that depends on the properties of the flow (see also figure 1). Note, that the relation (3) is consistent with the power-law scaling behavior of the energy spectral density in the inertial range, as discussed below (see equation (4)).
Within this rationale, the non-Gaussianity of the velocity increment distributions can be quantified by a single scalar parameter. The value of $\mu$ has been commonly accepted to be constant at large values of $Re_{\lambda}$. The reported values range between $0.2$ and $0.4$ for HIT and a value of $0.26$ has been established as representative of turbulent flows arneodo1996structure; vinnes2023characterizing; neunaber2020distinct.
Despite intense research on turbulent flows, no relation between $\mu$ and $C_{\varepsilon}$ has been predicted nor found. Furthermore, the universality of these constants remains an open question, with a growing amount of evidence pointing against this feature. While most studies deal with homogeneous turbulence, in this study we also focus on inhomogeneous or non-ideal turbulent flows. The data selection is explained below. To analyse such non-ideal and/or non-HIT data, we use equation (1) for $C_{\varepsilon}$ and equations 2 and 3 for $\mu$. For $C_{\varepsilon}$, we justify our approach by considering each velocity time series as a 1 D-surrogate. This was done before, for example for fractal-grid turbulence vassilicos2015dissipation and turbulent boundary layer flows nedic2017dissipation. For the estimation of $\mu$, we justify this approach based on the quality of the fits of the probability functions, $p(u_{r}(x))$, as shown in figure 9 in the appendix SM. Our approach aims to extend our understanding of ideal turbulent flow conditions. In particular, we aim to quantify the intermittency of such flows in a general way, independently of the turbulence model used to interpret the results. Therefore, a good fit of $p(u_{r}(x))$ is sufficient for our objectives. It is important to mention that the refined estimation of $\mu$ used in this work does not require an assumption regarding the behavior of the structure functions.
Hence, we present a detailed experimental work on turbulent wakes supported by previously published datasets on grid-generated turbulence and a turbulent axisymmetric jet.
Figure 1: Schematic representation of the spatial expansion of the turbulence cascade: the blue curve shows the energy spectrum, with energy transported to smaller scales, while the red curve illustrates the evolution of the shape factor $\Lambda^{2}(r)$ for normalized velocity increments. Note that in this sketch, $\Lambda_{0}^{2}$ is approximately zero.
I Experimental setup
We conducted wind tunnel experiments in LEGI, Grenoble by generating a turbulent wake behind porous and impermeable objects under different inflow conditions. The facility is a closed-circuit system with test section dimensions of 0.75 $\times$ 0.75 $\times$ 44\, m3 (see the appendix SM). The objects were cylinders and disks while both the diameter and solidity vary (see also table I in the appendix SM). To reduce the blockage, the cylinders were directly screwed to the walls of the tunnel while the disks were hold at the tunnel’s centreline with four thin piano wires which went through taped holes in the tunnel walls. For varying the inflow, both the entrance of the test section and the tunnel freestream velocity were modified. We used laminar inflow, and different turbulent background flows generated by a static regular grid with cylindrical bars or an active grid driven in a triple random mode (information about the turbulent flow produced by grids is reported in a previous work ferran2023experimental). The wake of the objects was scanned by a 11\,D-hot-wire with a temporal resolution of 5050\,kHz (with an anti-aliasing filter set at 3030\,kHz) which was traversed in streamwise and spanwise directions. The conversion from the temporal to the spatial domain was made using Taylor’s hypothesis. The nearest streamwise position to the wake generator was 5d5\,d and the farthest was 200d200\,d, with $d$ being the diameter of the object. For every position the data was acquired for 120120\,s. In total, $15$ configurations were carried out and $498$ points were taken. For simplicity, points in the neighborhood of the centreline are considered, including regions with stronger shear (pre-analysis shows that our results also hold for non-centreline data).
Additionally, the background flows without the object were also measured (see table I in the appendix SM for further information.)
As discussed above, we complement our study with $110$ velocity time series from three previously published experimental datasets covering different canonical flows ferran2023characterising; mora2019energy; renner2001experimental (for further details see the appendix SM). Within those measurement configurations, five involve decaying grid-generated turbulence and one captures a free axisymmetric jet. All cases use air as the fluid and the velocity time signals were collected by means of a 11\,D-hot-wire.
II Results
For all those points from all mentioned datasets, a rigorous analysis of common 11\,D-free-shear turbulence constants is derived and checked individually. Only time series that have the following characteristics were kept for further analysis:
a) the turbulence intensity $TI=u^{\prime}/\overline{u}$ with $\overline{u}$ being the actual mean velocity, decreases in streamwise direction, b) the large scale increments have an almost Gaussian distribution ($\Lambda_{0}^{2}<0.025$) c) the slope of the energy spectrum within the inertial range is smaller than 1.25-1.25, d) the length of the inertial range for the energy spectrum is larger than $0.5$ decades, e) the $R^{2}$-value of the linear regression of the fit of $\Lambda^{2}$ is larger than $0.99$, f) the mean squared deviation from the calculation of $\Lambda^{2}$ is in average smaller than $0.11$, g) the flatness of the velocity time signal is around 33 ($2.67<{\color[rgb]{0,0,0}F_{u}}<3.1$) and h) the skewness of the time signal is below $0.1$. Details and an example of the analysis is shown and discussed in the appendix SM. After this conditioning, 312 velocity time series remained for the analysis. Each of them represents at least $2\times 10^{4}$ integral time scales.
The integral length scale $L$ is calculated by integrating the autocorrelation function until it crosses $1/\mathrm{e}$ ($\mathrm{e}$ being the Eulerian number). Other integration limits were tested, and the trends presented in this Letter remained unchanged. $\varepsilon$ is derived – the common definition for HIT conditions is used – by integrating the dissipation spectrum $\varepsilon=\int_{0}^{\infty}15\nu\>k^{2}E(k)\>\mathrm{d}k$ (with $k$ the wavenumber in the Fourier space) and modeling the high-frequency noise with an extrapolated fit (a test with different fitting limits and functions gave an estimated error in the estimation of $\varepsilon$ in the order of $5\,\%$). We extracted, after equation (3), $\mu$ by fitting the evolution of the shape factor $\Lambda^{2}$ within the inertial range castaing1990velocity.
First, we investigate if the parameters $\mu$ and $C_{\varepsilon}$ present a constant value in our dataset, including all the flows studied. In figure 2a) it is clear that $\mu$ presents a large spreading and it does not approach towards an asymptotic limit at large values of $Re_{\lambda}$. (For reference, the dependency of $\mu$ with the streamwise distance to the turbulence generator is shown in the appendix SM). Figure 2b) shows that $C_{\varepsilon}$ also behaves similarly, i.e. $C_{\varepsilon}$ does not have a constant value.
Having a closer look on the data we see that for each set of boundary conditions, a linear relation between $C_{\varepsilon}$ and $\sqrt{Re_{G}}/Re_{\lambda}$ is consistent with the dissipation scalings from non-equilibrium turbulence vassilicos2015dissipation. Remarkably, when such conditions are changed, the linear relation still applies but with a different slope, showing the emergence of new length scales that affect the dissipation scalings.
Overall, our dataset does not show a universal or predictable behaviour in terms of neither $C_{\varepsilon}$ nor $\mu$. Next, we want to see if there is a relation between $\mu$ and $C_{\varepsilon}$. Figure 3a) shows that $\mu\propto 1/C_{\varepsilon}$, or, respectively, that the product of those two quantities is constant. In order to quantify this, initially, a simple fit is presented through the data points. Note that the commonly accepted value of $\mu$ of around 0.26 is not the asymptotic limit of our measurements. Moreover, the product of $\mu$ and $C_{\varepsilon}$, here defined as $\alpha$, is shown as a function of $Re_{\lambda}$ in figure 3b). Remarkably, $\alpha$ not only remains approximately constant for all cases, it also does not show any trends with $Re_{\lambda}$. Moreover, the well-centered PDF of $\alpha$ values indicates a Gaussian distribution. In the appendix we provide further figures and a table showing that the value of the standard deviation of $\alpha$ remains lower than the individual corresponding values of the standard deviation of $C_{\varepsilon}$ and $\mu$ SM. To the authors best knowledge, this is the first time a clear relation between these constants is found.
(a)                             (b)
Figure 2: a) $\mu$ as a function of $Re_{\lambda}$ for the highly restricted data. The red line corresponds to a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure. b) $C_{\varepsilon}$ versus $\sqrt{Re_{G}}/Re_{\lambda}$ for the highly restricted data. In general, the symbols in the legend are identical for all figures throughout this manuscript and correspond to the configurations in table I in the appendix SM. For laminar inflow, squared markers are used. For the regular grid and the active grid, markers are shaped as circles and triangles, respectively. For cylinders as generators, the markers contains a “//” (case names start with “C”) while for disks a $\backslash$” (case names start with “D”) is used. Hollow circular markers means no object and is equivalent to grid turbulence (case names start with “G”). The “x” marker indicates a free jet (case names start with “J”). $N_{v}$ indicates the number of velocity time series shown in this plot.
(a)                             (b)
Figure 3: a) $\mu$ as a function of $C_{\varepsilon}$ for the highly restricted data. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure, and the black line corresponds to a least-square fit ($R^{2}=0.86$ with the fitted values of $0.106\pm 0.002$ and $0.022\pm 0.009$). b) $\alpha$ versus $Re_{\lambda}$ for the highly restricted data with the PDF of $\alpha$ values $p(\alpha)$. The red line indicates the result for $\alpha$ from the fit from a) while the black solid line represents the actual mean value of the ensemble of $\alpha$ values. Additionally, two black dashed lines indicate the corresponding standard deviations $\sigma$ from the mean. For both figures, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the appendix SM. $N_{v}$ indicates the number of velocity time series shown in this plot.
We continue by investigating how two other relevant parameters behave, namely the Kolmogorov constant $C_{k}$ and the power law governing the energy spectral density within the inertial range. Following the notation of mydlarski1996onset,
$\color[rgb]{0,0,0}{E(k)=C_{k}\>\>\varepsilon^{2/3}\>k^{-5/3}\>(k\>\eta/2\>\pi)^{-\gamma+5/3},}$} (4)
the K41 prediction for the spectra is complemented with the correction $(k\>\eta/2\>\pi)^{-\gamma+5/3}$. K41 is therefore recovered when $\gamma=5/3$ and, for large values of $Re_{\lambda}$, $C_{k}$ is a constant (commonly accepted to be around 0.5 for HIT sreenivasan1995universality). The value of 5/3 emerges through a dimensional argument which followed the first two similarity hypotheses of Kolmogorov. Over the last years, several works reported values of $\gamma$ that deviate from the expected 5/3 value steiros2022balanced; rodriguez2023not; neunaber2020distinct, even at large values of $Re_{\lambda}$, most likely due to intermittency corrections. On the other hand, less evidence of deviations from $C_{k}\,\approx\,0.5$ has been reported in the literature meyers2008functional).
In figure 4a) & b), the dependency of $\gamma$ on $C_{\varepsilon}$ as well as $C_{k}$ on $\gamma$ are presented, respectively ($C_{k}$ as a function of $C_{\varepsilon}$ is shown in the appendix SM). Interestingly, the relation between $\mu$ and $C_{\varepsilon}$ is transferred to the other quantities. Moreover, all quantities present large variations with respect to the expected values for HIT within the K41 phenomenology. We remark that our results are still consistent with those reported in other experiments in inhomogeneous flows, such as other turbulent wakes neunaber2020distinct or atmospheric turbulence rodriguez2023not.
(a)                             (b)
Figure 4: a) $\gamma$ as a function of $C_{\varepsilon}$ and b) $C_{k}$ versus $\gamma$ for the highly restricted data. The red lines indicate both the commonly accepted value for $C_{k}$ and $\gamma$ for HIT sreenivasan1995universality, while the black lines correspond to a least-square fit for a) and b) with $R^{2}$ being $0.8$ and $0.99$, respectively. Once more, the symbols and corresponding configurations are shown and explained in figure 2 and table I in the appendix SM. $N_{v}$ indicates the number of velocity time series shown in this plot.
III Discussion
Based on the study of a very large number of velocity time series, the joint and individual behaviour of the dissipation constant and the intermittency factor were quantified, as well as the spectral law of turbulence. In contrast to the common approach of more restrictive data selection to obtain more precise results on turbulence, we adopted a less restrictive data selection method. This follows the idea of Vassilicos vassilicos2015dissipation, which allows for the investigation of states of non-equilibrium turbulence. For the different canonical turbulent flows that compose our dataset, our analysis leads to the proposition of a new relation,
$\mu\>C_{\varepsilon}=\alpha=\text{const.},$ (5)
replacing the previous proposition that these quantities should be individually constant.
Figure 3b indicates that there is no Reynolds number dependence of the product $\mu\>C_{\varepsilon}$. This is a remarkable result, which is additionally supported by the figures 12, 13 and 14 displayed in the appendix. It has long been claimed that certain turbulent properties are universal, i.e., independent of the type of flow, for large enough Reynolds numbers. Also, for free-shear anisotropic flows, a “tendency of recovery" of local isotropy is expected at high Reynolds numbers (see zhou2021turbulence). Together with the relations to other turbulence constants (shown in figure 4), this new proposed relation combines the dimensionless energy dissipation as an overall quantity of turbulent states with the statistics of velocity fluctuations at two points. Note that the power spectrum and the intermittency factor fully capture the two-point statistics morales2012characterization. Thus, we observe a new universality combining overall constants ($C_{\varepsilon}$ or $C_{k}$) with the two-point statistics of the velocity field.
As a remark, it should be pointed out that our results do not imply the absence of Reynolds number effects. For example, the depth of the cascade is still expected to depend on $Re_{\lambda}$, influencing the range of the power law of the power spectrum and therefore the intermittency at the smallest scales (which are also expected to be determined by the Reynolds number). We also observe that the evolution of $C_{\varepsilon}$ with the Reynolds number follows different linear relations sections for different flows, as discussed above.
Next, we set our result in the context of the $\Pi$-theorem of Vaschy-Buckingham vaschy1892lois; buckingham1914physically. $C_{\varepsilon}$ and $\mu$ can also be viewed as two $\Pi$-parameters for dissipation and the scale dependency of the fluctuations of dissipation. For different turbulent flow configurations, it is a challenge to find the functional relation between these $\Pi$-parameters. For a prominent example, see the Rayleigh-Bénard convection ahlers2009heat. In this sense, our result reads as $\mu\>C_{\varepsilon}=\Pi_{1}\cdot\Pi_{2}=\text{const.}\neq{\color[rgb]{0,0,0}\mathrm{func}(Re)}$.
Further, a relevant consequence of equation (5) is that we can define $\mu$ as,
$\mu=\frac{u^{\prime 3}\>\alpha}{\varepsilon\>L}.$ (6)
This relation has the advantage that the quantities on the right-hand side are much easier to measure and converge than $\mu$ (when estimated using standard approaches).
Next, we take a closer look at the quantities $\mu$ and $C_{\varepsilon}$ to understand the origin of the relationship given in equation (5). By re-writing equation (3) as,
$\color[rgb]{0,0,0}\mu\propto\frac{\mathrm{d\>\>}\Lambda^{2}(r)}{\mathrm{d\>\>ln}(r)},$} (7)
the intermittency factor $\mu$ can be interpreted as the “speed” of the cascade — that is, how rapidly the shape of the normalized velocity increment probability density function evolves from Gaussian behavior at large scales to heavy-tailed distributions at small scales. In contrast, the dissipation coefficient $C_{\varepsilon}$ can be re-expressed (introducing a large scale energy density rate $\varepsilon_{L}$) as,
$\color[rgb]{0,0,0}C_{\varepsilon}=\frac{\varepsilon}{\varepsilon_{L}}=\frac{\varepsilon}{\varepsilon_{L}(\varepsilon)},$} (8)
by combining $L$ and $u^{\prime 3}$, leading to $\varepsilon_{L}=u^{\prime 3}/L$. In this formulation, $C_{\varepsilon}$ in equation (8) represents the ratio between the actual dissipation rate $\varepsilon$ and $\varepsilon_{L}$, that acts as a surrogate of the input energy flux, associated with large-scale forcing. While $\varepsilon$ reflects the output of the cascade (the energy dissipated at small scales) $\varepsilon_{L}$ reflects the input energy at large scales. From an engineering point of view, $C_{\varepsilon}$ therefore quantifies the “efficiency” of the turbulence cascade in converting large-scale input into small-scale dissipation.
Given these interpretations, the observed relationship between $\mu$ and $C_{\varepsilon}$ is consistent with the second law of thermodynamics in out-of-equilibrium systems, which implies that faster processes tend to be less efficient callen1993thermodynamics. Turbulence, being a fundamentally non-equilibrium process, can be characterized thermodynamically by how far it deviates from equilibrium. We refer here to the thermodynamic equilibrium in the strict sense, i.e. the state of maximum entropy with no macroscopic fluxes, not to be confused with the equilibrium definition proposed in the context of “non-equilibrium turbulence”, which is discussed in the introduction. In this context, the values of $C_{\varepsilon}$ (and $\mu$) serve as a measure of this deviation. Figures 5 and 6 illustrate this behavior: as the system moves further from thermodynamic equilibrium, the cascade becomes faster (larger $\mu$), dissipation increases (larger $\varepsilon$), but efficiency decreases (smaller $C_{\varepsilon}$), and vice versa. Note that, unlike $C_{\varepsilon}$ and $\mu$ the variation of $\varepsilon$ considered here refers to comparisons to other measurement points within the same boundary conditions.
$\begin{array}[]{c}\text{fast process}\\ \mu\uparrow\end{array}\quad\widehat{=}\quad\begin{array}[]{c}\text{high dissipation}\\ \varepsilon\>{\color[rgb]{0,0,0}(\mathrm{case\;specific})}\uparrow\end{array}\quad\widehat{=}\quad\begin{array}[]{c}\text{low efficiency}\\ C_{\varepsilon}\downarrow\end{array}$
$\begin{array}[]{c}\text{slow process}\\ \mu\downarrow\end{array}\quad\widehat{=}\quad\begin{array}[]{c}\text{low dissipation}\\ \varepsilon\>{\color[rgb]{0,0,0}(\mathrm{case\;specific})}\downarrow\end{array}\quad\widehat{=}\quad\begin{array}[]{c}\text{high efficiency}\\ C_{\varepsilon}\uparrow\end{array}$
Figure 5: Relations of the intermittency factor $\mu$, the mean dissipation rate $\varepsilon$ and the dissipation constant $C_{\varepsilon}$.
Figure 6: Schematic representation of the relation between $\mu$ and $C_{\varepsilon}$. The yellow ellipses highlight the two boundary regions in the spectrum of $\mu$ and $C_{\varepsilon}$ values.
While our results appear to be applicable to a large variety of flows, further work is needed to assess whether our findings apply to other inhomogeneous and unsteady flows, which may lead to an even greater validity of our new law for turbulent flows.
Appendix A Further analysis of the dataset
In this appendix, we complement the main text showing a detailed example of our data analysis protocol. Additionally, further relations between constants are also shown and discussed, including a discussion about the validity of our results for our dataset processed with reduced quality contraints (in the sense of the restrictions discussed in section II).
Characterisation of data
In addition to the experimental dataset on turbulent wakes, which is explained in detail in the main manuscript, our study is extended by means of three other experimental studies: 1) decaying passive-grid generated turbulence within the Boundary Layer Wind Tunnel from the Laboratoire de Mécanique des Fluides in Lille (marked by G20, G21, G22). The measurement points are located between 3.3M3.3\,M and 54M54\,M, using $M$ as the mesh size. For further details, see ferran2023characterising. 2) Decaying passive-grid generated turbulence within the Lespinard Wind Tunnel from LEGI in Grenoble (marked by G23, G24). The measurement points are located between 1M1\,M and 30M30\,M, using $M$ as the mesh size. For further details, see mora2019energy. 3) An axisymmetric turbulent free jet (marked by J25). The measurement point is located at 125d125\,d, using $d$ as the nozzle diameter. For further details, see renner2001experimental.
For all added datasets, data acquisition was also achieved by using a 11\,D-hot-wire. The two studies investigating passive-grid turbulence were selected as one decays in agreement with Kolmogorov’s dissipation scalings kolmogorov1941local while the other study shows a decay with non-equilibrium behaviour vassilicos2015dissipation. The study of the jet was selected because it completes the series of canonical free-shear flows at steady state. Table 1 shows the entire dataset used in this paper and how different cases are marked as well as some further relevant details. Note that figure 7 completes the description of the dataset concerning turbulent wakes, detailed in the main manuscript, by showing a scheme of the experimental setup in LEGI, Grenoble.
Figure 7: Side view of the experimental setup in LEGI lab. The hot-wire can be traversed in the plane from the figure both in streamwise and spanwise directions (only centreline data is presented in this work). Both the grid and the object can be exchanged or removed individually.
$\textrm{L}_{1}$ $\textrm{W}_{1}$ $\textrm{W}_{2}$ $\textrm{T}_{2}$ ferran2023characterising mora2019energy renner2001experimental
[…]
Grid, $\overline{u}_{\infty}=10\,$m/s - G16 G17 - - - -
Grid, $\overline{u}_{\infty}=6\,$m/s - - G18 G19 - - -
Grid, $\overline{u}_{\infty}=7\,$m/s - - - - G20 - -
Grid, $\overline{u}_{\infty}=5\,$m/s - - - - G21 -
Grid, $\overline{u}_{\infty}=4\,$m/s - - - - G22 - -
Grid, $\overline{u}_{\infty}=8.6\,$m/s - - - - - G23 -
Grid, $\overline{u}_{\infty}=17\,$m/s - - - - - G24 -
Jet, $\overline{u}_{\infty}=45.5\,$m/s - - - - - - J25
[…]
Example of the analysis of one velocity time series
Figure 1 in the main manuscript presents a sketch that shows how the energy spectral density $E(k)$ and the shape factor $\Lambda^{2}(r)$ emerge across scales. In figure 8 we show a plot similar to figure 1 in the main manuscript extracted from a given velocity time signal from our wake experiment. The velocity time signal belongs to the case D15 from table 1. Both the energy spectral density and the shape factor have a clear power-law and logarithmic behaviour, respectively, over the inertial range. This is also indicated by the corresponding differentiation of this quantities which both present a plateau. In the figure, the scales used to fit both the energy spectral density and the shape factor are marked by differently hatched areas. Each of the velocity time series studied in this work was checked individually to verify whether the data and the fits were valid. Further details for this particular velocity time series are shown in table 2.
Figure 8: The evolvement of the energy spectral density $E(k)$ and the shape factor $\Lambda^{2}(r)$ are plotted over the scales $k$ and $r$, respectively. Additionally, the derivatives of the quantities in log/log and lin/log, respectively, are presented. Note that the axis correspond to the derivative of $\Lambda^{2}(r)$ already contains the pre-factor 9, so that the value of $\mu$ can be read directly. The fit areas are marked by hatched rectangles while the fits itself are visualised by solid black lines. Additionally, the value of $\Lambda^{2}=0$ is indicated with a black dashed line. Note that $\Lambda^{2}$ is almost 0 for very large increments.
$T$ (s) $f$ (kHz) $x/d$ $\overline{u}$ (m/s) $u^{\prime}/\overline{u}$ ($\%$) $Re_{\lambda}$ $S_{u}$ $F_{u}$ $L$ (cm) $\lambda$ ($m$) $\eta$ ($m$) $\varepsilon$ (m²/s³) $C_{\varepsilon}$ $\mu$ $\alpha$ ${\Lambda}^{2}_{0}$ $\gamma$ $C_{k}$
120 50 5 6.55 22.26 533 0.23 3.11 7.75 5.6 0.12 15.6 0.39 0.27 0.106 0.005 1.59 0.88
Table 2: The characteristics of the used velocity time signal in figure 8 are shown: $T$ is the sampling time, $f$ is the sampling frequency, $x/d$ is the streamwise distance normalized by the disk diameter, $\overline{u}$ is the mean velocity at the acquisition point, $u^{\prime}/\overline{u}$ is the turbulence intensity, $Re_{\lambda}$ is the Reynolds number based on the Taylor length scale, $S_{u}$ is the skewness of the velocity time signal, $F_{u}$ is the flatness of the velocity time signal, $L$ is the integral length scale, $\lambda$ is the Taylor length scale, $\eta$ is the Kolmogorov length scale, $\varepsilon$ is the mean dissipation rate, $C_{\varepsilon}$ is the dissipation constant, $\mu$ is the intermittency factor, $\alpha$ is the product of $\mu$ and $C_{\varepsilon}$, ${\Lambda_{0}}^{2}$ is the shape factor on the very large scales, $\gamma$ is the negative slope of the energy spectrum within the inertial range and $C_{k}$ is the Kolmogorov constant (according to equ. (4)).
Relation between $C_{k}$ and $C_{\varepsilon}$
Additionally to figure 4 in the main manuscript, the relation between $C_{k}$ as a function of $C_{\varepsilon}$ is presented for the highly restricted data.
Figure 9: $C_{k}$ as a function of $C_{\varepsilon}$ for the highly restricted data. The red line indicates the commonly accepted value for $C_{k}$ for HIT sreenivasan1995universality, while the black line corresponds to a least-square fit with $R^{2}$ being $0.79$. Once more, the symbols and corresponding configurations are shown and explained in figure 2 in the main manuscript and table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Relations between non-dimensional quantities with less restrictions
Figure 10 shows the relation between $\mu$ and $C_{\varepsilon}$ when the restrictions are softened with respect to the results presented in figure 3 from the main manuscript. In this case, the velocity time series that are considered, only have to fulfill the following conditions: a) $TI$ decreases in streamwise direction, b) the large scale increments have an almost Gaussian distribution ($\Lambda_{0}^{2}<0.025$) and c) the slope of the energy spectrum within the inertial range is smaller than 1.25-1.25, or in other words, is sufficiently steep. After this conditioning, 466 velocity time series remained for the analysis. For the less restricted data, the acquisition time represents at least $8\times 10^{3}$ integral time scales. It can be seen that the overall relation stays the same while the spreading becomes larger. Figure 11 shows the relation between the Kolmogorov constant $C_{k}$ and the exponent of the energy spectra $\gamma$ when the restrictions are softened in the same manner as mentioned before. It can be seen as well that the overall relation stays robust while the scatter becomes slightly larger.
Figure 10: $\mu$ over $C_{\varepsilon}$ for the less restricted dataset. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure, while the black line corresponds to a least-square fit ($R^{2}=0.77$). The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Figure 11: $C_{k}$ over $\gamma$ for the less restricted dataset. The red line indicates a commonly accepted values for $C_{k}$ and $\gamma$ for homogeneous isotropic turbulence sreenivasan1995universality, while the black line corresponds to a least-square fit ($R^{2}=0.99$) The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Independence of results on the Taylor-length based Reynolds number $Re_{\lambda}$
Figure 12 shows $Re_{\lambda}$ for each velocity time signal as a color map over the relation between $\mu$ and $C_{\varepsilon}$. It can be seen that the value of $Re_{\lambda}$ is not unique for neither $\mu$ nor $C_{\varepsilon}$. This becomes even clearer when looking at figure 13, where the plotted data is extended by only applying low restrictions. However, regarding $\alpha$ (defined in equation 5 in the main manuscript), there is no dependency on $Re_{\lambda}$ at all. This can be seen in figure 14, where $\alpha$ is directly plotted over $Re_{\lambda}$ for the less restricted dataset. There is no common trend towards small or large values of $Re_{\lambda}$. To quantify this observation the fit results of the plots of $\mu$ over $C_{\varepsilon}$ as well as the mean and median for the values of $\alpha$ for both sorts of restrictions are shown in table 3. The values of the fit results, mean and median are similar across the entire table. Furthermore, the table shows the values of the standard deviation of $\alpha$, $C_{\varepsilon}$ and $\mu$ for both sorts of restrictions. It is clear that the spreading around the mean is lower for $\alpha$ than for $C_{\varepsilon}$ and $\mu$.
Figure 12: For all data points within the highly restricted dataset, $Re_{\lambda}$ is presented by a color plot on top of the relation between $C_{\varepsilon}$ and $\mu$. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure while the black line corresponds to a least-square fit. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Figure 13: For all data points within the less restricted dataset, $Re_{\lambda}$ is presented by a color plot on top of the relation between $C_{\varepsilon}$ and $\mu$. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure, while the black line corresponds to a least-square fit. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Figure 14: For all data points within the less restricted dataset, $\alpha$ is presented over $Re_{\lambda}$ with the PDF of $\alpha$ values $p(\alpha)$. The red line indicates the result for $\alpha$ from the fit in figure 10. The black solid line represents the actual mean value of all plotted $\alpha$ values. Additionally, two black dashed lines indicate the corresponding standard deviations $\sigma$ from the mean. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
restriction fit result mean ($\alpha$) median ($\alpha$) $\sigma_{\alpha}$ $\sigma_{C_{\varepsilon}}$ $\sigma_{\mu}$
High 0.106 0.099 0.098 0.015 0.113 0.151
Low 0.105 0.1 0.097 0.021 0.117 0.152
Table 3: For both sorts of restrictions, the fit results of the plots of $\mu$ over $C_{\varepsilon}$ as well as the mean, median and standard deviation $\sigma$ for the values of $\alpha$ are shown. Additionally, the standard deviation $\sigma$ of the values of $\mu$ over $C_{\varepsilon}$ are presented.
Independence of results with the error within the calculation of the shape factor $\Lambda^{2}(r)$
As mentioned, the estimation of $\mu$ was validated for each velocity time series. Therefore, for each velocity time signal, two empirical probability density functions (PDF) of velocity increments for different scales $r$ were evaluated. To ensure the significance, all bins with less than 10 entries were removed. The chosen scales were the integral length scale $L$ and the Taylor length scale $\lambda$. The mean squared deviations between each empirical PDF of the velocity increments and the corresponding Castaing-curve $\hat{p}(u_{r}(x)/\sigma_{r})$} were calculated, obtaining $e_{c}(L)$} and $e_{c}(\lambda)$}. The overall deviation per velocity time series $\overline{e_{c}}$ is then calculated by computing the average of $e_{c}(L)$} and $e_{c}(\lambda)$}. The Castaing-curve can be obtained by using the equations 9 and 10 castaing1990velocity; morales2012characterization; neunaber2019stochastic. Figure 15 shows this method for one exemplary velocity time series. Figure 16 presents $\overline{e_{c}}$ for each velocity time series as a color plot over the relation between $\mu$ and $C_{\varepsilon}$. It can be seen that there are no trends with respect to this quantity.
${p(u_{r}(x)/\sigma_{r})}=$ (9)
$\frac{1}{2\pi\Lambda(r)}\>\int_{0}^{\infty}\>\frac{\mathrm{d}\sigma}{{\sigma}^{2}}\>\exp\left[-\frac{{u_{r}(x)}^{2}}{{2\sigma}^{2}}\right]\>\exp\left[-\frac{\mathrm{\ln}^{2}\left({\sigma/\sigma_{0}}\right)}{2\Lambda^{2}(r)}\right]$
$\sigma_{0}^{2}=\overline{u_{r}(x)^{2}}\>\exp\left[-2\Lambda^{2}(r)\right]$ (10)
Figure 15: Exemplary probability density functions of velocity increments for one velocity time series within the highly restricted dataset. In blue, the velocity increments for the integral length scale and in red for the Taylor length scale are presented. The shape factor at the integral length scale has a value of $0.043$ and at Taylor length scale a value of $0.125$. Additionally, for each of those two values of $r$, a Castaing-curve in black is plotted over the experimental data. For the Castaing-curve, the used value of $\Lambda^{2}$ corresponds to the extracted value from the experimental data. The mean squared deviation for the blue and red data from its Castaing-curve is $0.067$ and $0.069$, respectively. For reasons of visibility, the curves are shifted vertically.
Figure 16: For all data points within the highly restricted dataset, $\overline{e_{c}}$ is presented by a color plot on top of the relation between $C_{\varepsilon}$ and $\mu$. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure. The black line corresponds to a least-square fit. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
The distance between the object and the point of measurement
In figure 17, $\mu$ is shown over the non-dimensional streamwise distance to the turbulence-generating object from the perspective of the hot-wire probe. The streamwise distance $x$ is normalized by $d^{*}$, a typical characteristic length of the turbulence-generating object. For wakes, it is the diameter of the disk or cylinder while for grid-turbulence it is the mesh size and for the jet the nozzle diameter is used. For the shown highly restricted data, $x/d^{*}$ has a range between $4$ and $200$. It can be seen that high values of $\mu$ are not only a near-field phenomenon.
Figure 17: $\mu$ as a function of $x/d^{*}$, the non-dimensional streamwise distance to the turbulence generator for the highly restricted dataset. For the configurations with the disks and cylinders, $d^{*}$ is equal to the object diameter. For grid-generated turbulence, $d^{*}$ is equal to mesh size and for the jet, the diameter of the nozzle is taken. The red line indicates a commonly accepted value for $\mu$ for homogeneous isotropic turbulence arneodo1996structure. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
Distribution of skewness of velocity time signals
Figure 18 shows the distribution of the skewness $S_{u}$ of the velocity time series over $C_{\varepsilon}$ for the highly restricted dataset. Due to our data selection, where we are not restricting ourselves to HIT data, the values of $S_{u}$ are both positive and negative. Furthermore it can be seen that there is no clear trends with $C_{\varepsilon}$.
Figure 18: Skewness of the velocity time signal $S_{u}$ over $C_{\varepsilon}$ for the highly restricted dataset. The red line divides the data in positive and negative skewness. The symbols and corresponding configurations are shown in figure 2 from the main manuscript and explained in table 1. $N_{v}$ indicates the number of velocity time series shown in this plot.
ACKNOWLEDGMENTS
This work was supported by the LabEx Tec21 under Grant Investissements d’Avenir – grant agreement no. ANR-11-LABX-0030; and the Hanse-Wissenschaftskolleg (HWK Institute for Advanced Study, Delmenhorst, Germany) under a fellowship assigned to MO.
We thank Christophe Penisson, Muriel Lagauzere, Stephane Pioz-Marchand and Sylvain Dauge for helping with the experiment, Valentin Groß for proof-reading, Olivier De Marchi and Sebastian Bergemann for the IT support, Thomas Messmer, Lars Neuhaus and Daniela Moreno for validating parts of the used code and Michael Hölling, Ingrid Neunaber, Jan Friedrich and Christos Vassilicos for fruitful discussion.
DATA AVAILABILITY STATEMENT
The data that support the findings of this study are available from the corresponding author upon reasonable request.